Could you please help me with this problem. I am supposed to use a system of equations to solve this question.
I have 30 coins worth $6.00. The coins are quarters and dimes. How many of each do I have?
Thank you very much!

Answer by Alan3354(69443)

D + Q = 30 (30 Dimes and Quarters)
10D + 25Q = 600 ($6)
-------
Substitute for D (D=30-Q)
10(30-Q) + 25Q = 600
300 - 10Q + 25Q = 600
15Q = 300
Q = 20
D = 10
Check, it works.